4. The electrolytic cell contains 250.0 cm3 of an aqueous solution of 0,433 mol/l copper chloride CuCl2.

(a) For how long has the electrolysis been carried out with a current of 0,75 A when the analysis shows a copper (II) ion content of 0,167 mol/l after completion of the electrolysis?

b) How many grams of metallic copper is then formed on the cathode?

c) What reaction takes place at the anode? What's going on there?


5. Provide a chemical explanation for:

a) Iron nails are galvanized to prevent rusting.

b) Gold has been known since ancient times, iron was used at least as early as 2500-3000 BC, but the manufacture of aluminum objects has only begun since the end of the 19th century.

c) Write the chemical formulas of calcium compounds a - d.


+O2(g) +H₂O(l) +CO2(g) +HCl(aq)

a(s)---------→ a---------→ b---------→ c---------→ d


1
Expert's answer
2021-03-11T08:46:14-0500

Difference in concentration = (0.433-0.167)m/L = 0.266M

no of mole = volume × concentration

= 0.25L × 0.266M

= 0.0665mole


1 mole of Cu ion = 2 Faradays

= 2 ×96500C

0.0665mole = 2×96500×0.0665C

= 12,834.5C

Q=It

t= Q/I = 12834.5/0.75 = 17112.7 seconds


1 mole of copper =2F

64g of Cu = 2×96500C

2×96500C = 64g Cu

1C = 64/(2×96500C)

12834.5C =64 × 12834.5 / (2×96500)

= 4.256g


Oxidation takes place at the anode